# Sketch-Plan-Generalize: Continual Few-Shot Learning of Inductively Generalizable Spatial Concepts

### v2

@@ 6.3.2 Analysis of MCTS Variants for Concept Learning
Figure  4 compares the program accuracy for the three methods of plan search. For the third method the program accuracy is expected to be independent of expansion steps as it greedily chooses the action for which $\pi_{\text{neural}}$ gives the highest probability. For the MCTS+L based methods the accuracy increases beyond 0.6 with time, which demonstrates that having a composable library of concepts allows us to learn a much richer class of inductive concepts. MCTS+P+L saturates to a program accuracy of 0.933 in 4000 expansion steps, whereas MCTS-P+L takes 512000 expansion steps, which demonstrates a significant increase in learning efficiency via use of a neural prunner. For very low number of expansions steps (<40) accuracy of MCTS+L based methods is lower than MCTS-L as they spend some expansion steps on exploration based on UCB bound (rather than taking greedy choice).
6.3.3 Effectiveness of Continual Learning of Visual Concepts
Figure 5: Acquisition of new visual concepts. Likelihood of correct grounding of the referred object as training proceeds.
Having a disentangled representations allows us to 1) intersperse learning of new visual attributes with learning of new inductive concepts 2) avoid catastrophic forgetting of already learnt attributes. For example, the model can learn the chocolate from an instruction “construct a tower using chocolate blocks of size 4”, even if it has not seen the color in the pretraining phase. Because of our modular architecture, we can learn the color as a new embedding in the space of visual attributes. Plot  5 demonstrates the benefit of having such disentangled representations. As the training proceeds the probability of selecting the chocolate block increases with time. Further, the likelihood of selecting a magenta colored object (in a different scene) is unchanged and the problem of catastrophic forgetting is avoided. Additional details for continual learning of visual concepts appear in Appendix  D.2.
6.4 Application of Learnt Concepts for Embodied Instruction Following
Complex instruction execution via LLM. We demonstrate our ability to build over the learnt program representations and perform complex guided robot manipulation tasks, in a simulated Franka Panda environment. E.g., consider instructing a robot to : “Construct a tower of green die having the same height as the existing tower of white die.” or “Construct a tower of total 6 blocks using alternating blue and red blocks”. For both the above tasks we prompt GPT-4 by providing it with the set of inductive concepts that are already learnt, the set of primitive actions, and some pre-defined helper functions by using Python import statements in a manner similar to [code-as-policies]. GPT-4 generates an executable Python code in terms of these functions which on running generates the resultant trajectory. Top portion of figure  6 shows the execution for both these tasks. We provide details in Appendix  D.3.
Grounding learnt concepts into visual input for plan synthesis. We further demonstrate that the concepts we have acquired can help us to perform goal conditioned planning. Bottom portion of figure  6 demonstrates the results of our approach for the task of constructing a staircase beginning from an adversarial and assistive initial state. Note, the planner that we learn is a grounded neuro-symbolic planner, as a PDDL based planner cannot be hand-coded easily ( D.5), and LLMs/VLMs are unable to perform such complex reasoning tasks. Relevant details are provided in the Appendix  D.5
Figure 6: Application of learnt concepts. Top: Using LLM to generate the executable code for a novel tasks, given the concept definitions. Bottom: Integrating a neuro-symbolic planner over the concepts. In the bottom-left figure the planner optimally outputs the sequence of actions to replace the green cube. In the bottom-right, the planner is able to complete a staircase from an initially constructed row.
7 Limitations
Currently our methodology and evaluation setup assumes that each object is visible in the initial scene and does not handle partial observability. We assume all the objects are cubes of similar size in order to ensure dynamic stability of structures while stacking. During training we assume tracking and the ability to associate object with corresponding instances. We assume that the demonstration data given for learning is free of errors.
8 Conclusions
@@
This paper introduces a novel approach for learning inductive spatial concepts as neuro-symbolic programs from language-annotated demonstrations. We observe that direct generation of a program via an LLM/VLM does not generalize to new concepts and purely neural approaches fail to learn the inherent inductive prior inherent in such concepts. Our approach factors program learning as: Sketch: generating the high-level program signature via an LLM, Plan: searching for a grounded plan that maximises the total discounted reward with the respect to the demonstration, and Generalize: abstracting the grounded plan into an inductively generalize-able abstract plan via an LLM. Continual learning is achieved via learning of modular programs by giving preference to shorter programs (through composition of learnt ones) and allowing for learning of primitive neural concepts via gradient updates. Extensive evaluation demonstrates accurate program learning and stronger generalization in relation to baselines. Further, we show the grounding of learned concepts in visual data to facilitate reasoning and planning for embodied instruction following.
[…]
@@ Appendix A Additional Details on Technical Approach
Figure  7 illustrates the pipeline for learning inductive concepts.
Figure 7: SPG: Concept Learning Learning of grounded inductive programs is factored into three phases (a) Sketch: generating a sketch for the task intended in the instruction in the context of the environment (b) Plan: searching for the likely plan (guided by physical rewards) to construct the given assembly and (c) Generalize: generating a program representation to explain a class of demonstrations. Finally, our architecture allows learning of increasingly complex symbolic abstractions composed of simpler ones.
Figure  8 illustrates the pipeline for online inference to realize to realize construction of novel structures.

### v3

@@ 8 Results
Significance of MCTS in SPG. To assess the necessity and importance of MCTS, we carry out an ablation study where we replace it with an LLM planner during the planning stage, referred to as SPG-M+LMP. In the "plan" stage of our pipeline, GPT-4V is prompted to output a plan given the concept library and RGB keyframes from the demonstration. Our experiment shows that GPT-4V struggles to generate correct plans, particularly for complex structures like pyramids, arch_bridge and boundaries, resulting in significantly lower performance than SPG (Table 5). Additionally, some plans generated by GPT-4V are not physically grounded, leading to errors in both the planning and generalization stages, which compounds the inaccuracies. This demonstrates that combining symbolic search with LLMs offers a substantial advantage over using only LLMs.
Table 5: Ablation studies. Ablations with SPG-M+LMP and GPT-4V+VR. MSE values are in 1e-3.
Model Simple Complex
Acc. IoU MSE Acc. IoU MSE
SPG(Ours) 1.0 0.96 0.01 0.83 0.85 2.06
SPG-M+LMP 0.55 0.68 11.1 0.16 0.19 20.0
GPT-4V+VRF 0.66 0.75 6.8 0.16 0.46 12.0
Ablating with Pre-trained Models + Visual Reward Filter In line with program synthesis techniques using LLMs (Li et al., 2022; Chen et al., 2021), we sample five programs from GPT-4V and rank them according to the visual reward obtained from their execution. Furthermore, we provide the ground-truth programs of the concepts that are needed to learn the given new concept, thus employing a form of teacher forcing in program generation. Even with these measures, it performs significantly worse than SPG, see Table 5 (GPT-4V+VRF). While this performance is better than that of GPT-4V, it still unable to generate correct programs, especially for complex structures.
Q4: Incorporating Constraints into learnt concepts
Instructions with Discrete Constraints:Consider the instruction $\mathcal{\lambda}_{1}$: “Construct a staircase of size 5 such that all blocks have the same color as the block to their left. No block should have the same color as the block on top of it.” This example highlights the interpretive & compositional strength of our pipeline. We model the constraints as a CSP problem & use the Z3 solver (De Moura et al., 2008) to resolve them. The resulting grounded program for $\mathcal{\lambda}_{1}$ is executed via the planner, producing the desired 3D structure. Fig 6 shows the executed output in the simulator, satisfying both the structural concept and the imposed constraints.
Figure 6: Constructed staircase obeying all constraints in $\mathcal{\lambda}_{1}$, which are extracted via an LLM and solved via Z3 CSP-solver.
For simpler instructions such as $\mathcal{\lambda}_{2}$: ”Construct a tower of green die having the same height as the existing tower of white die.” and $\mathcal{\lambda}_{3}:$”Construct a tower of total 6 blocks using alternating blue and red blocks.”, the LLM (GPT-4) is able to understand the structural requirements (through the library $\mathcal{L}$) as well as the (simpler) constraints. Therefore, the pipeline (with access to tower through $\mathcal{L}$ or via SPG framework), generates Python code in terms of these functions, which on execution generates the required action sequence as shown in Fig. 7. (also see app.  D.3).
Figure 7: Using an LLM to solve the constraints in a novel task and generate executable code, given the concept definitions.
Efficient Plan Synthesis for Grounded Goals: As detailed in sec. 6 and App. D.5 , we demonstrate the ability to perform goal-conditioned planning using the acquired concepts. Fig 8 demonstrates the results of our approach for the tasks of constructing a staircase beginning from adversarial and assistive initial states.
Figure 8: Integrating a neuro-symbolic planner over the concepts. Top: The planner is able to optimally replace the green cube from the adversarial initial state by unstacking and re-stacking the faulty tower. Bottom: The planner is able to complete a staircase from an initially constructed row by layering rows upon rows, a method of construction it has not seen while learning staircase.
9 Conclusion
@@
This paper introduces a novel approach for learning inductive representation of grounded spatial concepts as neuro-symbolic programs via language-guided demonstrations. Our approach factors program learning as: Sketch: generating the high-level program signature via an LLM, Plan: searching for a grounded plan that maximises the total discounted reward with the respect to the demonstration, and Generalize: abstracting the grounded plan into an inductively generalize-able abstract plan via an LLM. Continual learning is achieved via learning of modular programs by giving preference to shorter programs through composition of learnt ones. Extensive evaluation demonstrates accurate program learning and stronger generalization in relation to purely LLM based as well as purely neural baselines. Grounding of learned concepts in visual data facilitates reasoning and planning for embodied instruction following. Limitations include reliance on perfect demonstrations, assumption of full observability of all objects and experiments confined to simulation. Incorporating noisy demonstrations, reasoning with beliefs and interleaving planning and execution remains part of future work.
@@ Appendix A Additional Details on Technical Approach
Figure  9 illustrates the pipeline for online inference to realize to realize construction of novel structures.